with explanation please question 7 it is with replacement QUESTION 5 Answer the question. Assume that a probability distribution is described by the discrete random variable x that ca
Answer #1

Question 7

The procedure results in binomial experiment because

1. there are 2 outcome.

a.) Red and P(Red)=12/40

b.) Non-red and P(Non-red)=28/40

2. Trials are independent.(with replacement)

3.) p=P(Red) remains same in each trial.